OpenStudy (anonymous):

Let Y1 and Y2 have the joint probability distribution function given by f(y1,y2)={4y1y2 0<=y1<=y2<=2 and 0 else. what is the joint distribution function?

OpenStudy (zarkon):

did you write the distribution down correctly?

OpenStudy (anonymous):

yes.

OpenStudy (zarkon):

when you integrate over what you have you get 8 (not 1)

OpenStudy (zarkon):

\[\int\limits_{0}^{2}\int\limits_{0}^{y_2}4y_1y_2\,dy_1dy_2=8\]

OpenStudy (anonymous):

okay, i must have calculated k wrong. when i calculated k I got 4

OpenStudy (zarkon):

ah

OpenStudy (anonymous):

Can you help me correctly calculate k?

OpenStudy (zarkon):

divide both sides by 8

OpenStudy (zarkon):

\[\int\limits_{0}^{2}\int\limits_{0}^{y_2}4y_1y_2\,dy_1dy_2=8\] \[\int\limits_{0}^{2}\int\limits_{0}^{y_2}\frac{1}{2}y_1y_2\,dy_1dy_2=1\]

OpenStudy (anonymous):

im sorry, the restrictions should be 0 to 1 not 0 to 2

OpenStudy (zarkon):

ah...ok

OpenStudy (zarkon):

still not 4

OpenStudy (anonymous):

is it then correct? or did I still incorrectly calculate k?

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\int\limits_{0}^{y_2}8y_1y_2\,dy_1dy_2=1\]

OpenStudy (anonymous):

can you show me how you got 8?

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\int\limits_{0}^{y_2}ky_1y_2\,dy_1dy_2\] \[=\int\limits_{0}^{1}\left[k\frac{y_1^2}{2}y_2\right]_{0}^{y_2}dy_2\] \[=\int\limits_{0}^{1}k\frac{y_2^2}{2}y_2dy_2\] \[=\int\limits_{0}^{1}k\frac{y_2^3}{2}dy_2\] \[=\left.k\frac{y_2^4}{8}\right|_{0}^{1}=\frac{k}{8}\] set \[\frac{k}{8}=1\] \[k=8\]

OpenStudy (anonymous):

ah i did my second integral wrong. Thank you.

OpenStudy (anonymous):

now back to my original question: joint distribution function

OpenStudy (zarkon):

\[F(y_1,y_2)=P(Y_1\le y_1,Y_2\le y_2)\]